Under standard conditions, what is the maximum electrical work, in joules, that the cell can accomplish if 85.0g of Sn is consumed? Answer Wmax in Joules.

A voltaic cell is based on the reaction :
Sn(s) + I2 (2) --> Sn^2+(aq) + 2I^-(aq)
Answer #1

1. convert grams to moles using molar mass of tin (118.7 g/mol) will give you .716 mole of zinc

2. set up the half reactions of the cell and determine the number of electrons transferred per mole of zn

3. determine the cell potential from the half reaction

4. dimensional analysis

.716mole Zn x (2mole e-/1mole Zn) x (96500 C/mole e-) x (.67V) = 9.26 x 10^4 J

-J= 1C x V

answered by: Danasia
Answer #2
To answer your question, I have to get the following Standard Reduction Potentials:
Sn(2+) + 2?e- ==> Sn(s), E° = -0.13V . . . (1)
I2(s) + 2?e- ==> 2?I-, E° = +0.54V . . . . . . (2)
(2)-(1):
Sn(s) + I2(s) ==> Sn(2+) + 2?I-, E° = 0.67V

Since the molar mass of Sn is: 118.7 g/mol, 85.0g Sn is:
(85.0g)/(118.7 g/mol) = 0.690 mole, the maximum electrical work, in joules, that the cell can accomplish is:
0.69*2*96485 = 130 (kJ)
answered by: Francisco Cantu
Answer #3
So one way to define a volt that helps us is:

1V = 1J/C (joule/coulomb)

To determine the voltage of this cell, we need to look up the standard reduction potentials for each half-reaction. They are:

Sn(2+)(aq) + 2e- --> Sn(s) Eo = -0.14 V
I2 (s) + 2 e- --> 2I- (aq) Eo = +0.54 V

Since we must have a final equation that cancels out the electrons and yields a positive voltage, we reverse the first equation and add up the Eo to give+0.68 V for the equation given initially. Also from the half-reactions we can see that for every mole of Sn consumed, 2 moles of electrons are transferredto iodine that can do work. To get the absolute number of electrons transferred, we need to determine how many moles of Sn are found in 85.0 g of Sn usingits molecular weight:

85.0g Sn / (78.96 g/mol) = 1.08 moles Sn consumed

Then multiply by 2 to get 2.16 moles of electrons transferred

To get the charge (in coulombs) of 2.16 moles of electrons, we use the Faraday constant, which is 96,485.3399 C/mol electrons. Thus:

(2.16 mol e-)*(96,485.3399 C/mole e-) = 208,408.3342 C

The maximum amount of work is then:

1 V = 1 J/C
1 J = 1 V*C
(0.68 V)*(208,408.3342 C) = 142 kJ
answered by: Demeisa